OpenStudy (anonymous):

Help please will give medal! Thomas Needs to cut a piece of wood for a model airplane he is building.He has four pices of wood:5 feet,4 feet,6 feet,and 3 feet long.if the lenth of wood he needs to cut is square root 11 feet long,which piece of wood should he cut to have the least amount of unused wood

OpenStudy (anonymous):

Remember, \(\sqrt9 = 3\) and \(\sqrt{16} = 4\), so \(\sqrt{11}\) must be in between 3 and 4.

OpenStudy (anonymous):

Thomas cannot cut \(\sqrt{11}\) with 3ft wood, because \(\sqrt{11}\) is greater than 3 so that left us 4ft, 5ft, and 6ft. which should he use to waste wood the least

OpenStudy (anonymous):

5ft?

OpenStudy (anonymous):

why not 4?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):